# InstructDET: Diversifying Referring Object Detection with Generalized Instructions

### v1

@@ Appendix D Visual Prompting and Visual-Textual Matching
Figure 12: A detailed pipeline of visual prompting and visual-textual matching. We use visual prompting to emphasize target object in the image. Then, we use the CLIP model to compute the scores for expression dropout.
Motivated by recent applications (Yang et al., 2023a; Shtedritski et al., 2023) of visual language pretraining (VLP) models in solving zero-shot REC tasks, we discover that coupling VLP with visual prompting enables robust pairing of image regions and text on the basis of potent generalization. The efficacy of pairing is largely influenced by the configuration of visual prompting, which can be divided into two types as follows:
@@
Once we obtain image $x$ and visually prompted $\widetilde{X}$, we send both of them to the CLIP model together with the generated expressions $Y$ from single modality and multi-modality pathways. Then, we can compute the global score $S_{g}$ and local score $S_{l}$, and thus obtain the final score $S_{f}$ as illustrated in equation 1 for expression dropout.
@@ Appendix E Instruction Generation for Mining Clusters of Multi-objects
[colback=gray!10,colframe=black,width=14cm, arc=3mm, auto outer arc, ] Task Description
You are an AI language assistant that can analyze phrases and sentenses. User will give you descriptions of several objects in an image. Descriptions of each object are several phrases or short sentences.
The given objects are expected to have similar properties. Based on the descriptions, find the common properties between given objects and summerize precisely as an assistant: common properties between objects can include same types, same functions, same color components, same poses, same relationships with other objects, engaging in the same activity, etc.
If there are no common properties between given objects, just tell that there are no common properties. Your summery should also be phrases. Do not repeat.
Give similarity between all given objects, contrary properties like different positions or different colors of clothes should not be included together in your descriptions.
One In-context Example
Prompt: Objects and their descriptions:
## object 2: girl sitting on bed, girl with toy, girl sitting on bed
## object 3: man looking down, boy sitting on the bed, man sitting on bed
Please find an summarize the similar properties of given objects.
Response: Summary of common properties of given objects:
## people on bed; person sitting on bed; people playing on bed; who sitting on bed;
@@ Appendix F Instruction Grouping
[colback=gray!10,colframe=black,width=14cm, arc=3mm, auto outer arc, ] Task Description
You are an AI language assistant that can analyze the language complexity of sentences or phrases.
User will give you a phrase or sentence describing a specific object in a image, which could be composed of nouns, adjectives, verbs, etc.
Grade the description according to its language complexity as an AI language assistant.
The language complexity of a phrase or sentence describing a specific object in an image can be graded into four levels:
Level 0. A single noun is used to give the object’s name or type.
Level 1. A phrase with one or more nouns, verbs and abjectives is used to describe simple attributes of the object itself, such as its color, its function or purpose, location in the image, or actions.
Level 2 A phrase with one or more nouns, verbs and abjectives is used to describe the object by referring to other objects in the image and describing their relationship, such as their relative positions or interactions.
Level 3. A long phrase or sentence is used to describe attributes of the object and also refer to a few other objects in detail, or describe a complicated or comprehensive/implicit relationship between multiple objects.
The level of descriptions increase as the language complexity and length increase, and also increase as the phrases or sentences become more descriptive and use more abjectives and nouns to describe the object.
One In-context Example
Prompt: Grade description: people who are sitting under an umbrella.
Response: My grading for description people who are sitting under an umbrella: This phrase is referring to the object of people, and gives simple object action of sitting and object relationship with the umbrella. The level of this description is: level 2.
@@ Appendix H Ablation Studies
In this section, we study the necessity of finetuning LLaVA during expression generation from multi-modality pathway. Then, we investigate how different visual prompting strategies affect the CLIP model to maintain our expected expressions.
@@
In our InstructDET, we do not finetune LLaMA in single-modality pathway, but finetune LLaVA in multi-modality pathway. There are two main reasons illustrated as follows:
@@
The selection of visual prompting (VP) is crucial for expression dropout. We evaluate the affect of various visual prompting methods based on the retrieval performance of CLIP model in Table 12. First, we define a new metric called expression retrieval ratio, which indicates the proportion of correct expressions that can be retrieved based on the visual-textual matching of CLIP model. During testing, each minibatch contains $k$ correct target expressions, and the remaining expressions are negative samples. We take the expressions with the top $k$ Clip score as the retrieved expressions. Finally, the proportion of correct expressions in the retrieved expressions is the expression retrieval ratio. Under the Easy setting, the negative samples in the minibatch come from other images. Under the Hard setting, the negative samples in the minibatch may come from different targets in the same image. Table 12 shows that cropping (1) and grayscale reversion (2) methods achieve poor results. Because cropping loses all surrounding environment information, and grayscale reversion loses all color features. The best single prompt is the contour blur reversion. The best combination prompt is the contour blur reversion and circle line. Circle line can indicate the rough areas that needs attention, and contour blur reversion can highlight the target object in a fine-grained manner that eliminates background interference. Inevitably, the prior knowledge in the CLIP model is also important to the difference in visual prompt effects. In CLIP’s pre-trained web-scale dataset, images with red circles often indicate that the targets in the red circles are more important and need to be noticed. A large amount of photography also exists in the web-scale dataset, which employs “Bokeh” to blur the background and highlight the subject.

### v5

@@ Appendix D Visual Prompting and Visual-Textual Matching
Figure 12: A detailed pipeline of visual prompting and visual-textual matching. We use visual prompting to emphasize target object in the image. Then, we use the CLIP model to compute the scores for expression filtering
Figure 12: A detailed pipeline of visual prompting and visual-textual matching. We use visual prompting to emphasize target object in the image. Then, we use the CLIP model to compute the scores for expression filtering
Motivated by recent applications (Yang et al., 2023; Shtedritski et al., 2023) of visual language pretraining (VLP) models in solving zero-shot REC tasks, we discover that coupling VLP with visual prompting enables robust pairing of image regions and text on the basis of potent generalization. The efficacy of pairing is largely influenced by the configuration of visual prompting, which can be divided into two types as follows:
@@
Once we obtain image $x$ and visually prompted $\widetilde{X}$, we send both of them to the CLIP model together with the generated expressions $Y$ from global prompt and local prompt pipelines. Then, we can compute the global score $S_{g}$ and local score $S_{l}$, and thus obtain the final score $S_{f}$ as illustrated in equation 1 for expression filtering.
@@ Appendix E Instruction Generation for Mining Clusters of Multi-objects
Task Description You are an AI language assistant that can analyze phrases and sentenses. User will give you descriptions of several objects in an image. Descriptions of each object are several phrases or short sentences. The given objects are expected to have similar properties. Based on the descriptions, find the common properties between given objects and summerize precisely as an assistant: common properties between objects can include same types, same functions, same color components, same poses, same relationships with other objects, engaging in the same activity, etc. If there are no common properties between given objects, just tell that there are no common properties. Your summery should also be phrases. Do not repeat. Give similarity between all given objects, contrary properties like different positions or different colors of clothes should not be included together in your descriptions. One In-context Example Prompt: Objects and their descriptions: ## object 2: girl sitting on bed, girl with toy, girl sitting on bed ## object 3: man looking down, boy sitting on the bed, man sitting on bed Please find an summarize the similar properties of given objects. Response: Summary of common properties of given objects: ## people on bed; person sitting on bed; people playing on bed; who sitting on bed;
@@ Appendix F Instruction Grouping
Task Description You are an AI language assistant that can analyze the language complexity of sentences or phrases. User will give you a phrase or sentence describing a specific object in a image, which could be composed of nouns, adjectives, verbs, etc. Grade the description according to its language complexity as an AI language assistant. The language complexity of a phrase or sentence describing a specific object in an image can be graded into four levels: Level 0. A single noun is used to give the object’s name or type. Level 1. A phrase with one or more nouns, verbs and abjectives is used to describe simple attributes of the object itself, such as its color, its function or purpose, location in the image, or actions. Level 2 A phrase with one or more nouns, verbs and abjectives is used to describe the object by referring to other objects in the image and describing their relationship, such as their relative positions or interactions. Level 3. A long phrase or sentence is used to describe attributes of the object and also refer to a few other objects in detail, or describe a complicated or comprehensive/implicit relationship between multiple objects. The level of descriptions increase as the language complexity and length increase, and also increase as the phrases or sentences become more descriptive and use more abjectives and nouns to describe the object. One In-context Example Prompt: Grade description: people who are sitting under an umbrella. Response: My grading for description people who are sitting under an umbrella: This phrase is referring to the object of people, and gives simple object action of sitting and object relationship with the umbrella. The level of this description is: level 2.
@@ Appendix H Ablation Studies
In this section, we study the necessity of finetuning LLaVA during expression generation from local prompt pipeline. Then, we investigate how different visual prompting strategies affect the CLIP model to maintain our expected expressions.
@@
In our InstructDET, we do not finetune LLaMA in global prompt pipeline, but finetune LLaVA in local prompt pipeline. There are two main reasons illustrated as follows:
@@
The selection of visual prompting (VP) is crucial for expression filtering. We evaluate the affect of various visual prompting methods based on the retrieval performance of CLIP model in Table 13. First, we define a new metric called expression retrieval ratio, which indicates the proportion of correct expressions that can be retrieved based on the visual-textual matching of CLIP model. During testing, each minibatch contains $k$ correct target expressions, and the remaining expressions are negative samples. We take the expressions with the top $k$ Clip score as the retrieved expressions. Finally, the proportion of correct expressions in the retrieved expressions is the expression retrieval ratio. Under the Easy setting, the negative samples in the minibatch come from other images. Under the Hard setting, the negative samples in the minibatch may come from different targets in the same image. Table 13 shows that cropping (1) and grayscale reversion (2) methods achieve poor results. Because cropping loses all surrounding environment information, and grayscale reversion loses all color features. The best single prompt is the contour blur reversion. The best combination prompt is the contour blur reversion and circle line. Circle line can indicate the rough areas that needs attention, and contour blur reversion can highlight the target object in a fine-grained manner that eliminates background interference. Inevitably, the prior knowledge in the CLIP model is also important to the difference in visual prompt effects. In CLIP’s pre-trained web-scale dataset, images with red circles often indicate that the targets in the red circles are more important and need to be noticed. A large amount of photography also exists in the web-scale dataset, which employs “Bokeh” to blur the background and highlight the subject.
Appendix I SYNONYMOUS REWRITE
In post processing, we utilize LLaMA to do synonymous rewriting to further diversify the generated expressions. The prompt we used is shown in Table 14.
Synonymous Rewriting Prompt I want you to act as a synonymous expression provider. I will give you a text of phrase or short sentence, which is an expression that describes a main object while mentioning some other objects. And you will reply to me with a new expression that have the same semantic meaning and describe the same main object as the provided expression. The new expressions should also be phrases or short sentences no longer than 25 words. Do not write explanations on replies. Do not repeat.
Table 14: Synonymous Rewriting Prompt for LLaMA.
Appendix J Supplement Comparisons
J.1 Evaluation on InDET
[…]
Method Backbone AP AP by Group
G1 G2 G3 G4 G5 G6
MDETR ResNet101 34.86 47.44 46.79 34.14 23.22 25.91 28.17
G-DINO SwinB 35.96 47.10 47.17 35.29 26.84 27.95 27.61
UNINEXT ResNet50 43.37 54.49 54.52 44.49 37.17 31.41 32.01
DROD (Ours)1 ResNet50 62.24 67.14 67.34 60.89 55.10 70.15 74.22
DROD (Ours)2 ResNet50 62.34 67.22 68.04 61.09 55.42 68.60 72.91
DROD (Ours) ViT-H 66.90 72.53 72.47 66.42 59.86 73.34 75.92
  • 1
    For fair comparison, our DROD model in Table 2 only utilizes RefCOCO/+/g datasets but with diversified instructions, which is partial of InDET.
  • 2
    Here we add evaluation results of DROD model trained on full InDET.
J.2 Evaluation on RefCOCO/g/+
[…]
Method Backbone Data RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val-u test-u
RefTR ResNet101 VG 85.65 88.73 81.16 77.55 82.26 68.99 79.25 80.01
MDETR ResNet101 GoldG,CC,RefC 86.75 89.58 81.41 79.52 84.09 70.62 81.64 80.89
PolyFormer SwinL GoldG,ReferIt,RefC 90.38 92.89 87.16 84.98 89.77 77.97 85.83 85.91
DROD (Ours)1 ResNet50 O365,CC,InDET 88.92 90.86 85.57 78.27 83.39 71.04 83.01 82.91
DROD (Ours)2 ResNet50 O365,CC,InDET 89.85 92.03 87.24 80.50 85.87 73.61 83.93 84.73
DROD (Ours)3 ViT-H O365,CC,InDET 92.93 94.47 91.13 86.20 89.82 80.86 88.62 89.46
  • 1
    For fair comparison, our DROD model in Table 3 only utilizes RefCOCO/+/g datasets but with diversified instructions, which is a part of InDET.
  • 2
    Here we add evaluation results of DROD model trained on full InDET.
  • 3
    DROD here with ViT-H as backbone also utilizes RefCOCO/+/g datasets with diversified instructions.
We have supplemented the comparisons with more specialized models, such as Polyformer(Liu et al., 2023b). With comparable magnitudes in the backbone network, our model continues to exhibit advantages.
Appendix K Time Consumption
Table 17: Time Consumption Statistics. We use ”instr.” as the short for ”instruction”.
Pipeline Step Main Model BatchSize Time Avg
Global Prompt Image Caption LLaVA 4 10.71 s/batch 2.68 s/image
Instr. Generation LLaMA 1 17.25 s/image 0.63 s/instr.
Local Prompt Instr. Generation LLaVA 16 2.47 s/batch 0.15 s/instr.
Post-Processing Instr. filtration CLIP 1 0.186 s/image 0.016 s/box
Multi-Object Instr. DBSCAN 1 0.053 s/image -
Instr. Grouping LLaMA 1 1.43e-06 s/instr. -
Table 17 shows the time consumption of the main steps in our instruction generation procedure. The data is obtained from NVIDIA-v100-32G machines. Obviously, the most time consuming steps are those who depend on large models. Except for the limited hardware performance, another reason why the instruction generation step in global prompt pipeline is very time consuming is that we have very long LLaMA prompts which include three in-context examples and we also desire long response which could include more diversified instruction for each object in image. So this step requires multiple machines to generate simultaneously. While the text received and generated in each forward pass is relatively short in the local prompt pipeline, the generation efficiency is significantly constrained by the fact that only one instruction can be generated per forward pass. Therefore, we address this limitation by maximizing parallelism through the increase in batch size, allowing the simultaneous generation of multiple instructions.